# LLMLight: Large Language Models as Traffic Signal Control Agents

### v4

@@ 4.2. Metrics
  • Average traveling time (ATT): The average traveling time quantifies the average duration of all the vehicles traveling from their origins to their respective destinations.
  • Average queue length (AQL): The average queue length is defined as the average number of queuing vehicles waiting in the road network.
  • Average waiting time (AWT): The average waiting time quantifies the average queuing time of vehicles at every intersection in the road network.
@@ 4.3. Compared Models
For transportation methods, we adopt Random, FixedTime, and Maxpressure as baselines. For RL methods, we compare MPLight, AttendLight, PressLight, CoLight, Efficient-CoLight, and Advanced-CoLight with our proposed method. For LLMs, we leverage GPT-4, ChatGPT-3.5, and the series of Llama-2 models as the backbone LLM (without fine-tuning) of LLMLight. Please refer to Appendix A.2 for more details about the abovementioned models.
@@ 4.4. Model Settings
All RL methods are trained with uniform hyperparameters, encompassing a learning rate of $1\times 10^{-3}$, a replay buffer size of $12000$, a sample size of $3000$, and a hidden size of $20$. The top_p value is standardized at $1.0$ for all LLMs. To stabilize the LLMs’ output, a temperature setting of zero is implemented for ChatGPT-3.5 and GPT-4, while Llama-2 is set to $0.1$. The LightGPT is developed from Llama-2-13B. In the imitation fine-tuning, the batch and mini-batch sizes are set to $128$ and $16$, respectively. We sample four reasoning trajectories under each instruction for critic-guided policy refinement. The fine-tuning learning rate is set to $3\times 10^{-4}$.
@@
Method Jinan Hangzhou
@@
LLMLight
Llama-2-13B 508.72 567.61 169.67 435.17 290.61 156.30 452.24 416.38 240.99 427.37 152.18 157.47 432.21 388.59 220.67
Llama-2-70B 353.03 286.52 85.96 324.52 162.34 99.87 320.41 210.13 100.90 357.95 93.21 106.63 361.53 250.69 121.58
@@
LightGPT (IFT) 277.60 166.17 48.21 272.24 104.93 48.78 261.34 129.27 45.02 316.4 61.55 49.92 333.42 190.43 68.99
LightGPT 274.03 159.39 43.24 266.94 100.46 40.34 260.17 127.08 41.00 311.76 57.89 42.84 330.71 189.09 64.16
@@ 4.5. Comparison Between the Traditional Methods and LLMLight (RQ1)
We first conduct experiments on Jinan and Hangzhou with five real-world traffic flow datasets. The comparison between traditional approaches and LLMLight is shown in Table 2. LightGPT (IFT) refers to the model trained exclusively through imitation fine-tuning. The experiment results reveal that LLMLight equipped with LightGPT consistently achieves state-of-the-art (SOTA) or comparable performance against all baselines. In detail, almost all transportation methods lag behind RL and LLMLight (LightGPT). Maxpressure demonstrates favorable performance in average waiting time (AWT) since it considers queuing vehicles the only privilege. The CoLight models exhibit the best performance among all RL methods, as they introduce neighboring information besides the target intersection. Although Advanced-CoLight, the SOTA RL method, outperforms LLMLight (LightGPT) on the Hangzhou datasets, it requires communication among neighboring intersections. On the contrary, LLMLight (LightGPT) demonstrates competitive results by leveraging observation features solely from the target intersection, underscoring its remarkable effectiveness.
As for LLMLight powered by generalist LLMs, we observe that GPT-4 outperforms all other LLMs and demonstrates competitive results against the most advanced RL methods, underscoring the effectiveness of our proposed prompting framework. Llama2-70B and 13B secure the second and third positions, indicating the adherence of LLMs to the scaling law (Kaplan et al., 2020) in the traffic signal control tasks. Surprisingly, ChatGPT-3.5 exhibits the least favorable performance. These experiment results indicate that generalist LLMs may lack expertise in traffic management, and there is a need to develop such domain-specific LLMs. The detailed analysis of their failure is elaborated in Appendix A.3.
@@
By fine-tuning to imitate the reasoning trajectories of GPT-4, LightGPT (IFT) demonstrates comparable or even better performance than GPT-4. This highlights the innate capability of the backbone model, Llama-2-13B, to excel in traffic signal control tasks if not constrained by its absence of traffic management expertise. Furthermore, The integration of policy refinement corrects the reasoning trajectories and decisions of the backbone LLM, resulting in a discernible enhancement. This implementation attains SOTA or competitive performance across all the datasets and significantly surpasses GPT-4, accentuating the promising prospects for advancing intelligent transportation-oriented LLMs.
@@ 4.6.1. Transferability
We first study the transferability ability of different methods by implementing the pre-trained model in another distinct road network. The experiment results are shown in Figure 4. Models labeled ”-T ” are pre-trained on a distinct road network (e.g., we evaluate the transferability in Hangzhou using the models pre-trained in Jinan). Otherwise, they are trained and tested in the same dataset. We observe a notable decline in the performance of reinforcement learning (RL)-based methods after the transfer, particularly evident in the Jinan 1 and Hangzhou 1 datasets. In contrast, LLMLight (LightGPT) consistently achieves superior performance and demonstrates remarkable transferability across all datasets. These findings highlight LightGPT’s exceptional generalization capability across diverse traffic contexts and robustness for practical implementations.
@@ 4.6.2. Scalability
Second, we delve into the scalability analysis of both traditional methods and LLMLight (LightGPT), scrutinizing their performance when applied to a considerably larger road network, New York, compared to the original road networks they are pre-trained on (i.e., Jinan and Hangzhou). The results of our experiments are detailed in Figure 5. We can observe significant performance degradations happening on most RL methods, as they demonstrate much worse performance than the heuristic method, Maxpressure. While the best RL methods exhibit suboptimal average travel time (ATT) in contrast to LLMLight (LightGPT), it is noteworthy that their policies result in significantly prolonged waiting times. Specifically, the waiting times are 57.80% and 49.80% longer than those achieved by ours. This observation suggests that the mechanism of the RL methods focuses on minimizing the overall number of queuing vehicles, possibly at the expense of causing extended waiting times for certain queues with fewer vehicles. The significance of minimizing waiting times cannot be overlooked in practical scenarios, as prolonged waits can induce anxiety and constitute a serious concern about the fairness of different drivers. In contrast, our proposed method ensures both the shortest travel and waiting times, highlighting its remarkable scalability and applicability when transitioning to much larger road networks.
@@ 4.6.3. Extreme High-traffic Scenarios
Next, we delve into an uncommon scenario when considerable traffic flow continuously appears at the intersection, a situation that rarely appears during training. To assess the efficacy of different methods in extreme high-traffic scenarios, we generate two synthetic traffic flow datasets on the Jinan and Hangzhou road networks, featuring an approximately fourfold increase in vehicles arriving within each 5-minute interval compared to the original datasets. Figure 6 shows the performance of both traditional methods and LLMLight armed with LightGPT (all the models are trained on Jinan 1 and Hangzhou 1 datasets, respectively). Similar to experiments on scalability, RL methods also showcase significant performance degradations, which exhibit poorer results than Maxpressure. This suggests that these pre-trained RL methods struggle to handle extreme high-traffic conditions when confronted with a significantly increased volume of vehicles compared to their training stage. In contrast, LLMLight (LightGPT) consistently shows superior performance, underscoring its robustness and practicality in much heavier traffic conditions.
@@ 4.7. Interpretability of LLMLight (RQ3)
To provide an in-depth assessment of LLMLight’s interpretability, we conducted a simulation case on the Hangzhou dataset. In Figure 7, we present the decision-making process of LightGPT. The detailed traffic conditions can be found in the Appendix A.3 (Table 3). In this simulated scenario, the north section experiences severe congestion, with a notable accumulation of queuing and approaching vehicles. LightGPT initiates its reasoning by comprehending the task and analyzing the traffic conditions at the targeted intersection. Subsequently, it adeptly identifies that the allowed lanes of signal NLSL demonstrate the heaviest congestion situation, thereby selecting it as the optimal signal phase to set. Unlike RL methods, LLMLight stands out not only for making effective control policies but also for its capacity to provide detailed explanations for each decision. This distinctive feature enhances the transparency and interpretability of LLMLight, contributing to a more informed understanding of its actions.
Figure 8. Prototype system.
4.8. Prototype System
We have successfully deployed a demonstration system designed to oversee the controlled traffic intersection and observe the decision-making of LightGPT. Figure 8 captures a snapshot of our demo system [^3] Our demo is available at https://llmtscs-demo.vercel.app/. In this setup, we provide users with real-time visualizations of the current traffic conditions on the right, along with statistical data of vehicles in the upper left corner. Additionally, the system further presents the reasoning behind LightGPT’s decisions through a message box in the interface’s lower left section.
@@ 5.1. Traffic Signal Control
Traffic signal control (TSC) methods can be broadly categorized into transportation (Koonce and Rodegerdts, 2008; Hunt et al., 1982) and reinforcement learning (RL)-based approaches (Chen et al., 2020; Wu et al., 2023). Initially, FixedTime (Koonce and Rodegerdts, 2008) employs a fixed cycle length and phase split to regulate traffic signals in a predetermined pattern. Maxpressure (Varaiya, 2013) greedily selects signal phases to maximize road network throughput. PressLight (Wei et al., 2019a) inherits the concept of Maxpressure and combines it with a RL algorithm to optimize the throughput of the intersection. Additionally, some studies propose to design more representative state features to describe traffic conditions. Advanced-MP (Zhang et al., 2022b) studies vehicle movement-level state features, designing Advanced-MP features based on traffic demand.
@@ 5.2. Intelligent Transportation
The application of intelligent transportation mainly focuses on forecasting and control. Within traffic flow forecasting (Liu et al., 2023, 2022; Ning et al., 2023), (Li et al., 2018) advocates for using graph neural networks (GNNs) to facilitate the exchange of information among proximate regions. Regarding control tasks, NavTL (Sun et al., 2023) presents an RL-based framework to manage traffic signal activation and autonomous vehicle rerouting in hybrid traffic scenarios. (Xu et al., 2022) developed a Double-Deep Q-Network (DDQN)-based model with fine-grained state features for efficient on-ramp traffic flow control. To facilitate sustainable transportation, (Zhang et al., 2022a) develop dynamic charging pricing for electric vehicles (EV) via a multi-agent graph convolutional RL model. (Lai et al., 2023) propose a meta-learning-based framework to provide personalized vehicle energy consumption predictions for various drivers.
@@ 5.3. Large Language Models for Decision Making
LLMs recently emerged and exhibited a remarkable capacity for zero-shot reasoning and generalization. Exploiting this proficiency, numerous studies (Fu et al., 2023; Cui et al., 2023; Chen et al., 2023b; Wen et al., 2023; Jin et al., 2023; Schick et al., 2023; Sha et al., 2023; Wang et al., 2023b; Da et al., 2024; Xu et al., 2023a) have increasingly employed LLMs as central planners in diverse control tasks. For instance, GLAM (Carta et al., 2023) trains the LLM control agents through interactive engagement with the environment. In intelligent transportation, DriveGPT4 (Xu et al., 2023b) presents an end-to-end autonomous driving system capable of interpreting vehicle actions and addressing diverse questions posed by human users. More studies on utilizing LLMs in intelligent transportation can be found at (Zhou et al., 2023; Zhang et al., 2024).
6. Conclusion and Findings
In this study, we introduce LLMLight, a novel framework that leverages large language models (LLMs) as traffic signal control agents. By instructing LLMs to conduct a human-like, step-by-step analysis of current traffic conditions, the intelligent agent could judiciously choose the optimal control action, thereby enhancing the overall traffic efficiency at the intersection. Furthermore, we propose a backbone LLM training procedure for optimizing a traffic signal control-oriented LLM, LightGPT. Through extensive experiments conducted on nine traffic flow datasets, the framework armed with LightGPT demonstrates remarkable improvements compared to traditional methods, underlying its exceptional effectiveness and generalization ability across diverse traffic scenarios.
Our findings are summarized as follows: (1) Traditional traffic signal control methods often exhibit limited generalization capabilities, leading to suboptimal transferability, scalability, and effectiveness in unconventional scenarios. (2) While general LLMs often lack a deep understanding of traffic management, employing a tailored prompting workflow and domain-specific fine-tuning has proven effective in overcoming these limitations and improving the generated control policies. (3) The LLM-empowered traffic signal control not only demonstrates superior performance across diverse traffic scenarios but also can offer detailed explanations for each decision. This would significantly contribute to a more comprehensible and accountable traffic control system.

### v5

@@ 4.2. Metrics
  • Average traveling time (ATT) quantifies the average duration of all the vehicles traveling from their origins to their respective destinations.
  • Average queue length (AQL) is defined as the average number of queuing vehicles waiting in the road network.
  • Average waiting time (AWT) quantifies the average queuing time of vehicles at every intersection in the road network.
@@ 4.3. Compared Models
For transportation methods, we adopt Random, FixedTime, and Maxpressure as baselines. For RL methods, we compare MPLight, AttendLight, PressLight, CoLight, Efficient-CoLight, and Advanced-CoLight with our proposed method. For LLMs, we utilize GPT-4, ChatGPT-3.5, Qwen, Llama2, and Llama3 as backbone models for LLMLight. For open-source LLMs, such as Qwen and Llama, we evaluate their performance both with and without fine-tuning. Due to the high training and deployment costs of models larger than 70B parameters (e.g., Llama2-70B or Qwen2-72B) in real-world traffic scenarios, we only test and report their performances without fine-tuning. Please refer to Appendix A.2 for more details about the abovementioned models.
@@ 4.4. Model Settings
All RL methods are trained with uniform hyperparameters, encompassing a learning rate of $1\times 10^{-3}$, a replay buffer size of $12000$, a sample size of $3000$, and a hidden size of $20$. The top_p value is standardized at $1.0$ for all LLMs. To stabilize the LLMs’ output, a temperature setting of zero is implemented for ChatGPT-3.5 and GPT-4, while open-sourced LLMs are set to $0.1$. The settings of LoRA, including a rank of 8 and a scaling factor $\alpha$ of 16, were set across all training stages. In the imitation fine-tuning, the batch and mini-batch sizes are set to $128$ and $16$, respectively. We sample four reasoning trajectories under each instruction for critic-guided policy refinement. The fine-tuning learning rate is set to $3\times 10^{-4}$.
@@
Models Jinan Hangzhou
@@
LLMLight (with Generalist LLMs)
Qwen2-72B 291.95 185.20 58.65 277.78 112.13 53.12 274.33 144.95 54.94 321.91 65.75 66.52 342.37 206.07 100.55
Llama2-70B 353.03 286.52 85.96 324.52 162.34 99.87 320.41 210.13 100.90 357.95 93.21 106.63 361.53 250.69 121.58
Llama3-70B 290.19 186.43 61.09 277.49 112.86 51.76 271.60 142.95 54.55 325.85 69.42 71.51 339.23 198.82 78.81
@@
LLMLight (with LightGPTs)
LightGPT (Qwen2-0.5B) 296.71 193.93 46.80 292.99 129.93 80.63 290.08 169.56 91.20 328.11 70.90 84.49 351.21 233.16 102.39
LightGPT (Qwen2-7B) 275.92 163.34 48.56 270.41 104.40 44.93 263.10 130.94 45.75 313.37 59.03 50.65 335.30 198.47 72.93
LightGPT (Llama2-7B) 275.11 161.35 46.38 269.01 102.92 43.06 260.53 127.75 41.84 314.24 59.59 39.66 333.94 191.63 65.49
LightGPT (Llama3-8B) 275.10 161.92 48.25 268.81 102.54 42.74 262.29 130.32 44.96 311.72 58.40 47.60 333.85 192.06 69.75
LightGPT (Llama2-13B) 274.03 159.39 43.24 266.94 100.46 40.34 260.17 127.08 41.00 310.78 56.93 38.64 330.71 189.09 64.16
@@ 4.5. Comparison Between the Traditional Methods and LLMLight (RQ1)
We first conduct experiments on Jinan and Hangzhou with five real-world traffic flow datasets. The comparison between traditional approaches and LLMLight is shown in Table 2. LightGPT refers to the backbone LLMs trained through our proposed fine-tuning method. The experimental results reveal that LLMLight, equipped with LightGPT, consistently achieves state-of-the-art (SOTA) or comparable performance against all baselines.
In detail, most transportation methods lag behind RL and LLMLight. Among RL models, the CoLight series stands out by integrating neighboring information beyond the target intersection. Although Advanced-CoLight, the SOTA RL method, slightly outperforms LLMLight on Hangzhou datasets, it requires communication among adjacent intersections. In contrast, LLMLight achieves competitive results using observation features solely from the target intersection, showcasing its notable effectiveness. To provide a fair comparison, we also compare Advanced-CoLight without the agent cooperation mechanism and LLMLight, detailed in Appendix A.10. The results show that LLMLight excels in all metrics under this fair setting. We believe LLMLight can be further improved in the future by incorporating multi-agent cooperation.
In our evaluation of LLMLight powered by generalist LLMs, GPT-4 excels and competes with top RL methods, highlighting the effectiveness of our proposed control framework. Llama3-70B and Qwen2-72B rank second and third. Surprisingly, ChatGPT-3.5 shows the least favorable performance, indicating that generalist LLMs may lack the specialized expertise for optimal traffic management. This suggests a potential need for domain-specific LLMs for intelligent traffic control in the future. Appendix A.8 provides more performance comparisons and analysis of their limitations.
Examining the models trained by our proposed method (i.e., LightGPTs), even LLMLight powered by the smallest model, LightGPT (Qwen2-0.5B), outperforms the much larger Llama2-70B. This highlights the effectiveness of our backbone model training approach. While lightweight models with 7B or 8B parameters exhibit a slight performance drop compared to LightGPT (Llama2-13B), they offer more economical alternatives for deployment. Please refer to Appendix A.4 for a detailed cost-effectiveness analysis.
Table 3. Abalation tests of four LightGPT variants.
Model Average Travel Time (ATT, s)
GPT-4 Qwen2-0.5B Qwen2-7B Llama2-7B Llama3-8B Llama2-13B
Origin 318.71 1124.84 621.80 973.50 1136.17 722.73
IFT (W/O filtering) - 788.93 314.33 316.43 316.70 313.32
IFT - 331.32 314.32 315.94 314.56 312.43
IFT + CGPR - 328.11 313.37 314.24 311.72 310.78
4.6. Abalation Study
To verify the effectiveness of each stage in our proposed backbone LLM training method, we conduct ablation studies on four variants of our proposed LightGPT: (1) Origin: The original LLM without any fine-tuning. (2) IFT (W/O filtering): The LLM trained only with imitation fine-tuning, excluding the high-quality reasoning trajectory filtering strategy. (3) IFT: The LLM trained only with imitation fine-tuning (IFT). (4) IFT + CGPR: The LLM trained through all stages, including imitation fine-tuning (IFT) and critic-guided policy refinement (CGPR). The experiment results on five backbone LLMs are reported in Table 3. The results demonstrate that the effectiveness of each backbone LLM significantly improves after applying our proposed fine-tuning strategy. Furthermore, each training stage consistently enhances the control policy of the LLMs, as evidenced by the progressively improved performance with the incorporation of each proposed stage.
@@ 4.6. Abalation Study
By fine-tuning to imitate the reasoning trajectories of GPT-4, LightGPT demonstrates comparable or even better performance than GPT-4. This highlights the innate capability of the backbone model to excel in traffic signal control tasks if not constrained by their absence of traffic management expertise. Furthermore, The integration of policy refinement corrects the reasoning trajectories and decisions of the backbone LLM, resulting in a discernible enhancement. This implementation attains SOTA or competitive performance across all the datasets and significantly surpasses GPT-4, accentuating the promising prospects for advancing intelligent transportation-oriented LLMs.
@@ 4.7. Generlization Comparision (RQ2)
In this section, we evaluate the generalization ability of traditional methods and our proposed LLMLight in unseen environments. LLMLight is powered by the LightGPT pre-trained on Llama2-13B. All models undergo uniform pre-training using either the Jinan 1 or Hangzhou 1 dataset. An additional 24-hour cycle experiment is detailed in Appendix A.5.
@@ 4.7.1. Transferability
We first study the transferability ability of different methods by implementing the pre-trained model in another distinct road network. The experiment results are shown in Figure 4. Models labeled ”-T ” are pre-trained on a distinct road network (e.g., we evaluate the transferability in Hangzhou using the models pre-trained in Jinan). Otherwise, they are trained and tested in the same dataset. We observe a notable decline in the performance of RL-based methods after the transfer, particularly evident in the Jinan 1 and Hangzhou 1 datasets. In contrast, LLMLight consistently achieves superior performance and demonstrates remarkable transferability across all datasets. These findings highlight LightGPT’s exceptional generalization capability across diverse traffic contexts and robustness for practical implementations.
@@ 4.7.2. Scalability
Second, we delve into the scalability analysis of both traditional methods and LLMLight, scrutinizing their performance when applied to a considerably larger road network, New York, compared to the original road networks they are pre-trained on (i.e., the Jinan 1 dataset). The results of our experiments are detailed in Figure 5. We can observe significant performance degradations happening on most RL methods, as they demonstrate much worse performance than the heuristic method, Maxpressure. While the best RL methods exhibit suboptimal average travel time (ATT) in contrast to LLMLight, it is noteworthy that their policies result in significantly prolonged waiting times. Specifically, the waiting times are 57.80% and 49.80% longer than those achieved by ours. This observation suggests that the mechanism of the RL methods focuses on minimizing the overall number of queuing vehicles, possibly at the expense of causing extended waiting times for certain queues with fewer vehicles. The significance of minimizing waiting times cannot be overlooked in practical scenarios, as prolonged waits can induce anxiety and constitute a serious concern about the fairness of different drivers. In contrast, our proposed method ensures both the shortest travel and waiting times, highlighting its remarkable scalability and applicability when transitioning to much larger road networks.
@@ 4.7.3. Extreme High-traffic Scenarios
Next, we delve into an uncommon scenario when considerable traffic flow continuously appears at the intersection, a situation that rarely appears during training. To assess the efficacy of different methods in extreme high-traffic scenarios, we generate two synthetic traffic flow datasets on the Jinan and Hangzhou road networks, featuring an approximately fourfold increase in vehicles arriving within each 5-minute interval compared to the original datasets. Figure 6 shows the performance of both traditional methods and LLMLight armed with LightGPT. All models are pre-trained on regular traffic flow datasets from the same road network (i.e., the Jinan 1 or Hangzhou 1 dataset). Similar to experiments on scalability, RL methods also showcase significant performance degradations, which exhibit poorer results than Maxpressure. This suggests that these pre-trained RL methods struggle with out-of-distribution (OOD) traffic conditions when confronted with a significantly increased volume of vehicles compared to their training stage. In contrast, LLMLight consistently shows superior performance, underscoring its robustness and practicality in much heavier traffic conditions.
@@ 4.8. Interpretability of LLMLight (RQ3)
To assess LLMLight’s interpretability, we analyze a scenario in the Hangzhou dataset. Figure 7 illustrates the decision-making process of LightGPT (Llama2-13B). Detailed traffic conditions are reported in Appendix A.8 (Table 7). In this scenario, The north section encounters severe congestion with a significant queue. LightGPT first understands the task and analyzes traffic at the intersection. Subsequently, it adeptly identifies that the allowed lanes of signal NLSL demonstrate the heaviest congestion situation, thereby selecting it as the optimal signal phase to set. Unlike RL methods, LLMLight not only excels in making effective control policies but also in providing detailed explanations for each decision. This unique feature enhances the transparency and interpretability of LLMLight, contributing to a more informed understanding of its actions.
4.9. Deployment Analysis
We are partnering with the Institute of Deep Perception Technology [^2] http://www.idpt-jitri.com/ in China to deploy LLMLight in real-world transportation systems [^3] A prototype system for the public is available at https://lightgpt2024.github.io/LLMLight_Demo/.. We conduct a human expert evaluation of LLMLight across five randomly selected traffic scenarios to assess its reasoning and decision-making capabilities in various traffic scenarios. Feedback was gathered from a diverse group: four traffic police officers from the Kunming Southeast Ring Expressway Jinsong Traffic Police Brigade in Yunnan and the Zhongshan Traffic Police Detachment in Guangdong, China, nine experienced drivers, and two AI specialists. The consensus from these evaluations was overwhelmingly positive, with all participants acknowledging LLMLight’s reasoning ability and decisions made in all examined scenarios. They highly praised its applicability and notable interpretability in real-world contexts. The detailed evaluation results, studied traffic scenarios, and LightGPT’s reasoning processes are shown in Appendix A.6.
Additionally: (1) We conduct a cost-effectiveness analysis comparing LightGPT with GPT-4, detailed in Appendix A.4, which includes training and deployment budgets and efficiency evaluations. (2) A 24-hour experiment in Appendix A.5 validated LightGPT’s robustness under daily long-term traffic fluctuations. (3) In Appendix A.7.1, we discuss potential challenges and data privacy preservation methods in real-world deployment. In conclusion, LightGPT consistently demonstrates superior cost-effectiveness and applicability for real-world deployment scenarios.
@@ 5.1. Traffic Signal Control
Traffic signal control (TSC) methods fall into two main categories: transportation methods and reinforcement learning (RL)-based approaches. Transportation methods include FixedTime (Koonce and Rodegerdts, 2008), which uses a set cycle length and phase split, and Maxpressure (Varaiya, 2013), which optimizes signal phases to maximize network throughput. PressLight (Wei et al., 2019a) extends Maxpressure by integrating RL to enhance intersection throughput. Recent studies also focus on improving state feature representation for better traffic condition modeling. For instance, Advanced-MP (Zhang et al., 2022b) develops advanced state features based on vehicle movement and traffic demand.
@@ 5.2. Intelligent Transportation
Intelligent transportation applications primarily emphasize forecasting (Liu et al., 2023, 2022; Ning et al., 2023) and control (Sun et al., 2023; Xu et al., 2022). For instance, (Lai et al., 2023) presents a meta-learning-based framework designed to deliver personalized vehicle energy consumption predictions tailored to individual drivers. Additionally, (Zhang et al., 2022a) introduces a dynamic charging pricing model for electric vehicles (EVs) using a multi-agent graph convolutional reinforcement learning approach.
@@ 5.3. Large Language Models for Decision Making
LLMs recently emerged and exhibited impressive zero-shot reasoning and generalization abilities. Exploiting this proficiency, numerous studies (Fu et al., 2023; Cui et al., 2023; Chen et al., 2023b; Wen et al., 2023; Jin et al., 2023; Schick et al., 2023; Sha et al., 2023; Wang et al., 2023b; Da et al., 2024; Xu et al., 2024; Feng et al., 2024a) have increasingly employed LLMs as central planners in various control tasks. For example, (Feng et al., 2024b) introduces CityBench, an interactive simulation-based benchmark for evaluating LLMs’ capabilities in urban domain perception and decision-making. DriveGPT4 (Xu et al., 2023) showcases an end-to-end autonomous driving system that interprets vehicle actions and responds to user queries. Additional studies on LLMs in intelligent transportation are detailed in (Zhou et al., 2023; Zhang et al., 2024).
6. Conclusion and Limitations
This study introduces LLMLight, a framework employing LLMs as traffic signal control agents. By performing a human-like, step-by-step analysis of traffic conditions, LLMLight adjusts traffic lights to improve traffic efficiency. We further propose a backbone LLM training procedure for developing the TSC-oriented LLM, LightGPT. Extensive experiments on ten traffic datasets show that LLMLight with LightGPT significantly outperforms nine baseline methods and ten leading LLMs, demonstrating its superior effectiveness, generalization, and interoperability across diverse traffic scenarios.
Our findings include: (1) Traditional traffic signal control methods have limited generalization capabilities, leading to suboptimal performance in unseen traffic scenarios. (2) Generalist LLMs lack in-depth traffic management understanding, but tailored prompting and domain-specific fine-tuning can significantly improve their control policies. (3) LLM-empowered traffic signal control demonstrates superior performance across diverse scenarios and can provide detailed explanations for each decision, contributing to a more comprehensible and accountable traffic control system.
Our limitations include: (1) Our approach focuses solely on traffic features at the target intersection, neglecting multi-agent cooperation mechanisms for global traffic improvement. (2) We do not incorporate additional modality features, such as camera images, to aid the LLM agent in decision-making. (3) The framework primarily regulates vehicle traffic flow and does not consider the impact of other dynamic factors at the intersection, such as pedestrians and bicycles. We plan to address these limitations to enhance the effectiveness of the LLMLight framework in the future.